# GAS: Generative Avatar Synthesis from a Single Image

### v1

@@ Abstract
We introduce a generalizable and unified framework to synthesize view-consistent and temporally coherent avatars from a single image, addressing the challenging problem of single-image avatar generation. While recent methods employ diffusion models conditioned on human templates like depth or normal maps, they often struggle to preserve appearance information due to the discrepancy between sparse driving signals and the actual human subject, resulting in multi-view and temporal inconsistencies. Our approach bridges this gap by combining the reconstruction power of regression-based 3D human reconstruction with the generative capabilities of a diffusion model. The dense driving signal from the initial reconstructed human provides comprehensive conditioning, ensuring high-quality synthesis faithful to the reference appearance and structure. Additionally, we propose a unified framework that enables the generalization learned from novel pose synthesis on in-the-wild videos to naturally transfer to novel view synthesis. Our video-based diffusion model enhances disentangled synthesis with high-quality view-consistent renderings for novel views and realistic non-rigid deformations in novel pose animation. Results demonstrate the superior generalization ability of our method across in-domain and out-of-domain in-the-wild datasets. Project page: https://humansensinglab.github.io/GAS/
@@
1 Introduction
Figure 1: In-the-wild avatar synthesis across views and time. Left: The avatar is animated based on a pose sequence, enabling novel view generation from any chosen synthesized frame (highlighted by a red dashed rectangle). Right: Starting from a reference image (highlighted by a red dashed rectangle), we generate its novel views and animate the avatar through a given pose sequence.
Human avatar generation is a longstanding research area in computer vision and graphics, with applications spanning gaming, film, sports, fashion, and telepresence. Despite its potential to revolutionize daily life, current avatar generation technologies require costly capturing setups and highly complicated procedures, making them hard to access by broader audiences.
Recent efforts have sought to make avatar generation more affordable by harnessing the power of neural rendering. In particular, some works [21, 22, 17, 23] focus on generalizable 3D human reconstruction that synthesize novel views and poses of arbitrary human subjects from very sparse or even single input images. These methods utilize 3D human priors to enable robust synthesis of complex human geometry, and interpolate well between available observations. However, due to their regression-based nature, they often struggle with extrapolation, leading to blurry results. In addition, they are often limited to rigid deformations.
@@ 1 Introduction
Generative models like GANs and diffusion models have recently shown remarkable capability in producing photorealistic detailed images and videos with realistic dynamics. Building on the success of these generative techniques, several approaches use diffusion models conditioned on human priors (e.g., 2D keypoints, depth maps, or normal maps) to create high-quality avatars with realistic animation from a single image [16, 63, 6, 41]. However, due to the sparsity in these conditioning signals, the generated avatars often suffer from inconsistencies such as flickering across different views and over time.
@@
To tackle these challenges, we propose a method for single-image avatar synthesis that achieves both view and temporal consistency. Instead of relying solely on sparse signals, we first generate intermediate novel views or poses using a regression-based 3D human reconstruction model. We then use them as conditioning inputs for a video diffusion model. By incorporating dense information from the 3D reconstruction, our approach preserves both structural accuracy and rich visual details, ensuring high-quality, consistent results across views and time.
@@
Beyond selecting the appropriate driving signal, another key challenge is improving generalization to real-world data, because ultimately we want a system that works not just on controlled datasets but also on casually captured in-the-wild human images. Most human digitization models are trained on multi-view datasets captured in controlled studios [21, 22, 17], which lack the diversity of real-world scenarios, such as varied lighting, clothing, and motion. To address this, we propose to incorporate in-the-wild internet videos, which provide an abundant source of diverse real-world distribution. While leveraging such videos significantly improves generalization, training a model on them presents challenges—particularly for novel view synthesis, which typically requires multi-view supervision, rarely available in internet videos.
@@
To overcome this, we propose a unified framework that jointly learns novel view and pose synthesis, sharing parameters across tasks to enable cross-task generalization. While we still use studio multi-view datasets for novel view synthesis, we incorporate both multi-view and in-the-wild videos for novel pose synthesis. Since the same model parameters are used for both tasks, the generalization gained from pose synthesis naturally transfers to novel view synthesis as well. This approach significantly enhances the model’s ability to handle real-world data and improves overall synthesis quality as can be seen in Figure 1. Additionally, in order to employ this unified framework effectively, we introduce a mode switcher that differentiates between novel view and pose synthesis. This switcher disentangles the tasks, allowing the network to focus on view consistency when synthesizing novel views and realistic deformations when generating novel poses.
@@
  • A dense appearance cue as synthesis guidance, ensuring consistent appearance preservation.
  • A switcher integrated into the video diffusion model to enable disentangled modeling of static novel views and dynamic pose animation.
@@ 2.1 Generalizable Human Radiance Fields
Radiance fields methods, like NeRF [31], have shown strong capabilities in generating high-fidelity renderings across views. To extend these methods to generalize to new scenes and handle sparse-view inputs, several approaches [54, 40, 45] have incorporated pixel-aligned features to aid novel view synthesis. However, applying these approaches directly to human modeling presents challenges due to the complex geometry of the human body, including self-occlusions, which often lead to overly smoothed outputs. To address these challenges, some methods [21, 22, 59, 34, 35, 8, 9] use 3D human templates, such as the SMPL model [30], to anchor features accurately on the human form. Although these techniques achieve good results even with sparse or single-view inputs [17], they are often hindered by slow rendering speeds. Recently, 3D Gaussian Splatting [20], optimized with GPU-based rasterization, has emerged as a promising radiance field representation, enabling faster rendering. Building on this, new methods [62, 23, 64] have achieved photorealistic human rendering from sparse observations. Nevertheless, these methods still struggle with producing clear details in unseen regions due to mean-seeking behavior in one-to-many mapping scenarios [28, 24]. In this work, we propose learning a distribution conditioned on generalizable human radiance fields, allowing us to synthesize clean, consistent view and pose sequences by sampling from this distribution.
@@ 2.2 Generative Human Animation
Generative human animation aims to employ generative models to produce coherent videos from static human images, utilizing guidance such as text and motion sequences. This body of work focuses on leveraging generative priors to sample complex dynamic motions, including pose-dependent clothing deformations. Early approaches harnessed the generative capabilities of Generative Adversarial Networks (GANs) [11] for synthesizing novel human poses [47, 5, 27]. In recent years, latent diffusion models [39] have gained traction in the realm of human animation due to their robust controllability and superior generation quality. Various methods [46, 6, 16, 63, 26, 42, 7, 25] implement distinct motion guidance and conditioning techniques. Notably, Animate Anyone [16] introduces a UNet-based ReferenceNet to extract features from reference images, utilizing DWPose [52] for pose guidance. Subsequent works [63] also incorporate guidance from 3D human parametric models, such as SMPL [30, 60, 61], leveraging the advantages of multiple forms of guidance. Following this trajectory, recent studies [14, 41, 36, 29] explore human view controllability within the diffusion frameworks. Human4DiT [41] develops a hierarchical 4D diffusion transformer that disentangles the learning of 2D images, viewpoints, and time. However, these methods struggle to synthesize view-consistent and temporally coherent results due to the gap between the sparse driving signal and the actual subject. In this paper, we propose a unified approach to human view and pose synthesis by conditioning on generalizable geometry and appearance features and mapping them to a cohesive representation, resulting in a more straightforward and effective process.
@@ 2.3 Diffusion Models for Video Generation
Diffusion models have demonstrated exceptional performance in image synthesis over the past few years. The research community has since started to extend these models to the more challenging task of video generation [56]. Significant efforts have been made to adapt UNet-based image diffusion models for video generation by incorporating additional temporal modules. Several studies [3, 12] have proposed freezing pre-trained image diffusion models and training only the newly added temporal module on video data. Stable Video Diffusion [2] adds temporal layers after each spatial convolution and attention layer, fine-tuning the entire model on large-scale, curated video datasets. It has proven to be a strong foundational model, facilitating further advancements in video generation as well as 3D/4D synthesis [44, 49]. Beyond UNet-based diffusion models, diffusion transformers [33, 53, 4] have emerged as a robust architecture for video generation, applying full attention to space-time patches of video and image latent codes. In this work, we opt to leverage the strong foundational capabilities of Stable Video Diffusion to model both multi-view correspondences and temporally coherent realistic deformations of human subjects.
@@ 3 Method
Figure 2: Overview of our proposed method. From a single input human image, our approach leverages a generalizable human NeRF to map the subject to canonical space, followed by reposing and rendering to obtain the appearance cue (i.e., NeRF renderings). Paired with the geometry cue (i.e., SMPL normal maps), they provide comprehensive conditions for the video diffusion model, enabling multi-view consistent novel view synthesis or temporally consistent pose animation, effectively disentangled by a switcher module.
Our method generates view and time-wise coherent avatar renderings from a single image by leveraging the dense appearance information of a generalizable human reconstruction model combined with the generative capabilities of a video diffusion model.
@@ 3.2 Single-view Generalizable Human Synthesis
The target image $I_{tar}$ in target view and pose is rendered using the volume rendering [31].
We emphasize that this human NeRF model is designed to generalize, enabling the synthesis of human appearances across a range of novel views and poses. However, due to its mean-seeking tendency, producing sharp renderings—particularly in occluded or unseen regions—remains challenging. This limitation motivates our introduction of generative priors to learn a distribution over the human views and poses, as discussed in the next section.
@@ 3.3 Synthesis as Video Generation Condition
In this section, we elaborate how we leverage video diffusion prior to reformulate the novel view and pose synthesis as a video generation task conditioned on the human radiance fields and human geometric templates. We would like to note that prior approaches often treat novel view and novel pose synthesis as separate tasks. In contrast, our approach unifies both tasks by conditioning on NeRF-rendered images, allowing for more consistent renderings. Using human templates alone as a condition provides only sparse geometric guidance, which often leads to inconsistencies in appearance across novel views. By integrating generalizable radiance field as a unified conditioning mechanism, we address this issue and achieve better appearance consistency across varied poses and views.
Novel View Synthesis. Given a reference image $I_{ref}$ and a camera trajectory $\{P_{1},P_{2},\cdots,P_{T}\}$, we render the corresponding images $\bm{I}_{nerf}=\{I_{1},I_{2},\cdots,I_{T}\}$ from the NeRF, which servers as an input to our video diffusion model, i.e., Stable Video Diffusion [2] that adopts spatio-temporal attention module and 3D residual convolution in the diffusion UNet. For the single input image $I_{\text{ref}}$, we extract its feature with CLIP [37] and repeat it for $T$ times to match the number of frames, denoted as $\bm{h}_{\text{clip}}$, which is then added to the video diffusion model through the cross attention. Meanwhile, we use the VAE encoder to encode our input image $I_{\text{ref}}$ and obtain its latent feature $C_{\text{vae}}$.
@@ 4.1.1 Datasets
3D Scans. We use THuman2.1 [55] and 2K2K dataset [13] with around 4500 3D scans in total. THuman2.1 comprises 2445 high-quality 3D scans and texture maps. We randomly sample 2345 subjects for training and the remaining 100 subjects for testing. 2K2K dataset consists of 2000 train and 50 test subjects, totally 2050 scans. For both datasets, RGB images are rendered from 20 uniformly distributed views around the scan, at the resolution of $1024\times 1024$. SMPL parameters are estimated using off-the-shelf method for multi-view SMPL fitting [1].
@@ 4.1.3 Baselines and Metrics
Figure 3: Qualitative comparisons for novel view synthesis on THuman dataset. Inconsistent areas across views are marked with red circles, e.g., the hair regions are inconsistent between frontal and back views. * denotes a method fine-tuned on our full 3D scan dataset.
Figure 4: Qualitative comparisons for novel pose synthesis on MVHumanNet. In the first row of Champ [63] results, blue rectangles mark disappearing arms, green rectangles show varying sleeve lengths, and red rectangles indicate inconsistencies in hair appearance. * denotes a method further fine-tuned using our complete animation video dataset.
@@ 4.2 Comparisons on Novel View Synthesis
We show the comparisons with Animate Anyone [16] and Champ [63], which are two competitors focusing on generative human synthesis. For a fair comparison, we fine-tuned both methods on our full 3D scan dataset for 10k iterations.
Quantitative results. We adopt a comprehensive evaluation protocol for human novel view synthesis from a single image. Specifically, for each subject, we sample four orthogonal input views and compute their mean score across all generated novel views. It is worth noting that Animate Anyone [16] often produces a noisy background. For a fair comparison that focuses solely on the human subjects, we apply the ground truth mask to remove backgrounds in the THuman dataset, while in the 2K2K dataset, we retain the generated images as they are.
@@
Qualitative results. Figure 3 presents our qualitative comparisons with baseline methods, focusing on challenging reference views and novel views that have less overlapping.
@@ 4.3 Comparisons on Novel Pose Synthesis
We compare our method on novel pose synthesis task with strong baseline methods Animate Anyone [16] and Champ [63], which are designed for single-image animation. For a fair comparison, we also further fine-tuned the baseline methods on our complete video dataset for 10k iterations.
@@
Qualitative results. Qualitative results are shown in Figure 4, where we compare our method with Champ [63] on the novel view animation task for the MVHumanNet dataset. Competing methods display flickering and fluctuations, especially in unseen regions. In contrast, ours synthesizes temporally consistent animations, even from novel views.

### v2

@@ Abstract
We present a unified and generalizable framework for synthesizing view-consistent and temporally coherent avatars from a single image, addressing the challenging task of single-image avatar generation. Existing diffusion-based methods often condition on sparse human templates (e.g., depth or normal maps), which leads to multi-view and temporal inconsistencies due to the mismatch between these signals and the true appearance of the subject. Our approach bridges this gap by combining the reconstruction power of regression-based 3D human reconstruction with the generative capabilities of a diffusion model. In a first step, an initial 3D reconstructed human through a generalized NeRF provides comprehensive conditioning, ensuring high-quality synthesis faithful to the reference appearance and structure. Subsequently, the derived geometry and appearance from the generalized NeRF serve as input to a video-based diffusion model. This strategic integration is pivotal for enforcing both multi-view and temporal consistency throughout the avatar’s generation. Empirical results underscore the superior generalization ability of our proposed method, demonstrating its effectiveness across diverse in-domain and out-of-domain in-the-wild datasets.
Figure 1: In-the-wild avatar synthesis across views and poses. Starting from a reference image, we generate its novel views and animate the avatar given a pose sequence.
1 Introduction
Human avatar generation has been a longstanding focus in computer vision and graphics, with wide-ranging applications in gaming, film, sports, fashion, and telepresence. However, despite its transformative potential, existing technologies often rely on expensive capture systems and complex workflows, making them inaccessible to the broader public. Recent advancements have aimed to make avatar generation more accessible by leveraging neural rendering techniques. In particular, studies such as [23, 24, 17, 25] explore generalizable 3D human reconstruction, enabling novel view and arbitrary pose synthesis of subjects from highly sparse—sometimes even single—input images. These approaches incorporate 3D human priors to support accurate synthesis of complex body geometry and smooth interpolation between input observations. However, their regression-based nature limits their extrapolation capabilities, often resulting in blurry outputs and restricting them to mostly rigid deformations.
@@ 1 Introduction
Generative models, such as GANs and diffusion models, have recently demonstrated impressive capabilities in synthesizing photorealistic images and videos with realistic motion. Leveraging these advancements, recent approaches employ diffusion models conditioned on human priors—such as 2D keypoints, depth, or normal maps—to generate high-quality avatars from a single image  [16, 67, 6, 44]. However, the sparsity of these conditioning cues often leads to inconsistencies in the generated results, including flickering across views and temporal instability.
@@
To address these challenges, we introduce a method for single-image avatar synthesis that ensures both view and temporal consistency. Rather than depending solely on sparse conditioning signals (e.g., SMPL normal maps), we first generate intermediate novel views and/or poses using a regression-based 3D human reconstruction model. The shape and apperance of these new views/poses are then used as inputs to a video diffusion model. By leveraging the dense structural information from the 3D reconstruction, our approach maintains geometric fidelity and visual richness, delivering high-quality, consistent results across viewpoints and human poses (i.e., time).
@@
Another major challenge lies in improving generalization to real-world data, as the ultimate goal is to develop a system that performs well beyond controlled environments—specifically on casually captured, in-the-wild human imagery. Most existing human digitization models are trained on multi-view datasets collected in studio settings [23, 24, 17], which lack the diversity found in real-world scenes, including variations in lighting, clothing, and movement. To overcome this limitation, we propose leveraging in-the-wild internet videos, which offer a rich and diverse distribution of real-world appearances. However, training with such data introduces new challenges, particularly for novel view synthesis, which traditionally relies on multi-view supervision—an element typically absent in internet-sourced footage.
@@
Our method, Generative Avatar Synthesis (GAS) is a unified framework that jointly learns both novel view and pose synthesis by sharing parameters across tasks. While studio-captured multi-view datasets are used for view synthesis, we augment training with both multi-view and in-the-wild videos for pose synthesis. This parameter sharing allows improvements from pose generalization to transfer naturally to view synthesis, significantly enhancing performance on real-world data (see Figure 1). To enable this joint learning effectively, we incorporate a mode switcher that distinguishes between the two tasks. This design allows the network to prioritize view consistency for novel view synthesis and realistic deformation for novel pose generation.
@@
  • A dense appearance cue derived from generalizable NeRF renderings as diffusion guidance, ensuring consistent appearance preservation over novel views and poses.
@@ 2.1 Generalizable Human Radiance Fields
Radiance field methods like NeRF [35] have demonstrated impressive performance in generating high-fidelity novel views. To extend these models to generalize across scenes and work with sparse-view inputs, several approaches [59, 43, 48] incorporate pixel-aligned features. However, applying them directly to human modeling remains challenging due to complex body geometry and self-occlusions, often resulting in over-smoothed outputs. To address this, recent works [23, 24, 63, 37, 38, 8, 9] leverage 3D human priors such as the SMPL model [34] to better anchor features on the human form, enabling robust synthesis from sparse or even single-view inputs [17]. However, these methods typically suffer from slow rendering speeds.
To overcome this, 3D Gaussian Splatting [22], accelerated by GPU rasterization, has emerged as an efficient radiance field representation. Recent works [66, 25, 68] leverage this framework for photorealistic human rendering from sparse inputs. Yet, they still face difficulties in generating fine details in unseen regions due to the mean-seeking nature of one-to-many mappings [32, 26]. Moreover, animating avatars typically requires costly post-processing—e.g., using linear blend skinning [27]—which often fails to capture non-rigid effects such as garment dynamics [68, 17, 23, 39, 18, 51].
In this work, we propose a novel framework that learns a distribution shift over generalizable human radiance fields, enabling consistent and artifact-free synthesis of novel views and poses. By incorporating a strong generative prior, our method naturally produces realistic non-rigid deformations from pose sequences without relying on complex post-processing pipelines.
@@ 2.2 Generative Human Animation
Generative human animation aims to employ generative models to produce coherent videos from static human images, utilizing guidance such as text and motion sequences. This body of work focuses on leveraging generative priors to sample complex dynamic motions, including pose-dependent clothing deformations. Early approaches harnessed the generative capabilities of Generative Adversarial Networks (GANs) [11] for synthesizing novel human poses [50, 5, 31]. In recent years, latent diffusion models [42] have gained traction in the realm of human animation due to their robust controllability and superior generation quality. Various methods [49, 6, 16, 67, 29, 45, 7, 28] implement distinct motion guidance and conditioning techniques. Notably, Animate Anyone [16] introduces a UNet-based ReferenceNet to extract features from reference images, utilizing DWPose [57] for pose guidance. Subsequent works [67] also incorporate guidance from 3D human parametric models, such as SMPL [34, 64, 65], leveraging the advantages of multiple forms of guidance. Following this trajectory, recent studies [14, 44, 39, 33, 21, 56, 30] explore human view controllability within the diffusion frameworks. Human4DiT [44] develops a hierarchical 4D diffusion transformer that disentangles the learning of 2D images, viewpoints, and time. However, these methods struggle to synthesize view-consistent and temporally coherent results due to the gap between the sparse driving signal and the actual subject. In this paper, we address this challenge by densely conditioning on generalizable geometry and appearance cues, leading to improved appearance preservation and consistency.
@@ 2.3 Diffusion Models for Video Generation
Diffusion models have shown impressive results in image synthesis and are now being adapted to the more complex domain of video generation [61]. A common strategy involves extending UNet-based image diffusion models by adding temporal modules. Some approaches [3, 12] freeze the pre-trained image backbone and train only the temporal components on video data. Stable Video Diffusion (SVD) [2], for instance, enhances the UNet architecture by inserting temporal layers after each spatial convolution and attention block, and fine-tunes the full model on large-scale curated video datasets. This approach has proven to be a powerful foundation, enabling downstream applications such as video generation and 3D/4D synthesis [47, 53]. In parallel, diffusion transformers [36, 58, 4] have emerged as an alternative architecture, applying full spatio-temporal attention over latent codes for both images and videos. In this work, we build on the strong capabilities of SVD to model multi-view consistency and temporally coherent, realistic deformations of human subjects.
@@ 3 Method
Figure 2: Overview of GAS. Starting from a single input image, GAS uses a generalizable human NeRF to map the subject into a canonical space, then reposes and renders the 3D NeRF model to extract detailed appearance cues (i.e., NeRF renderings). These are paired with geometry cues (i.e., SMPL normal maps) and fed into a video diffusion model. A switcher module disentangles the tasks, enabling the model to generate either multi-view consistent novel views or temporally coherent pose animations.
Figure 2 illustrates the main idea of our method. GAS synthesizes view- and pose-consistent avatar renderings from a single image by combining the rich appearance information from a generalizable 3D human reconstruction model with the generative power of a video diffusion model.
@@ 3.2 Single-view Generalizable Human Synthesis
The target image $I_{tar}$ , corresponding to the desired view and pose, is generated using volume rendering as described in [35]. We emphasize that this human NeRF model is designed to generalize, enabling the synthesis of human appearances across a range of novel views and poses. However, due to its mean-seeking property, producing sharp renderings, particularly in occluded or unseen regions, remains challenging. This limitation motivates our introduction of generative priors to learn a distribution over the human views and poses, as discussed in the next section.
@@ 3.3 Synthesis as Video Generation Condition
This section describes how we leverage a video diffusion prior to reformulate novel view and pose synthesis as a unified video generation task, conditioned on human radiance fields and geometric templates. While previous methods typically treat view and pose synthesis as separate problems, our approach bridges them by conditioning the diffusion model on NeRF-rendered images. This dense and appearance-rich signal leads to more consistent outputs. In contrast, conditioning solely on sparse human templates, as in prior work [67], often results in inconsistent appearance across views.
Novel View Synthesis. Given a reference image $I_{ref}$ and a camera trajectory $\{P_{1},P_{2},\cdots,P_{T}\}$, we render the corresponding images $\bm{I}_{nerf}=\{I_{1},I_{2},\cdots,I_{T}\}$ from the NeRF, which servers as an input to our video diffusion model, Stable Video Diffusion (SVD) [2]. SVD has a spatio-temporal attention module and 3D residual convolution in the diffusion UNet. For the single input image $I_{\text{ref}}$, we extract its feature with CLIP [40] and repeat it for $T$ times, denoted as $\bm{h}_{\text{clip}}$, which is then added to the video diffusion model through the cross attention. Meanwhile, we use the VAE encoder to encode our input image $I_{\text{ref}}$ and obtain its latent feature $C_{\text{vae}}$.
@@ 4.1.1 Datasets
3D Scans. We use THuman2.1 [60] and 2K2K dataset [13] with around 4500 3D scans in total. THuman2.1 comprises 2445 high-quality 3D scans and texture maps. We randomly sample 2345 subjects for training and the remaining 100 subjects for testing. 2K2K dataset consists of 2000 train and 50 test subjects, totally 2050 scans. For both datasets, RGB images are rendered from 20 uniformly distributed views around the scan, at the resolution of $\times 1024$ 1024. SMPL parameters are estimated using off-the-shelf method for multi-view SMPL fitting [1].
@@ 4.1.3 Baselines and Metrics
Figure 3: Qualitative comparisons for novel view synthesis on the THuman dataset. For the first subject, our method generates cleaner garment textures and sharper facial details, achieving better realism and consistency across views (e.g., the hair style in our generated front view is faithful to the reference image). For the second subject, baseline methods exhibit inconsistencies across views, marked with circles (e.g., hair misalignment between generated front and back views). * denotes fine-tuning on our full 3D scan dataset.
Figure 4: Qualitative comparisons for novel pose synthesis on MVHumanNet dataset. In the first row of Champ [67] results, blue rectangles mark disappearing arms, green rectangles show varying sleeve lengths, and red rectangles indicate inconsistencies in hair appearance. * denotes a method further fine-tuned using our complete animation video dataset.
@@ 4.2 Comparisons on Novel View Synthesis
We compare our approach with two leading generative human synthesis methods—Animate Anyone [16] and Champ [67]. For a fair evaluation, we fine-tuned both models on our complete 3D scan dataset for 10,000 iterations.
Quantitative results. We employ a rigorous evaluation protocol for single-image human novel view synthesis. For each subject, we sample four orthogonal input views and report the average performance across all corresponding novel views. Notably, Animate Anyone [16] tends to generate noisy backgrounds. To ensure a fair comparison that emphasizes the synthesized human subject, we apply the ground truth masks to remove backgrounds in the THuman dataset. In contrast, for the 2K2K dataset, we evaluate the generated outputs without background removal.
@@
Qualitative results. Figure 3 presents our qualitative comparisons with baseline methods, focusing on challenging reference views and novel views with less overlap.
@@ 4.3 Comparisons on Novel Pose Synthesis
We compare GAS on novel pose synthesis task with Animate Anyone [16] and Champ [67]. For a fair comparison, we also further fine-tuned the baseline methods on our complete video dataset for 10k iterations.
@@
Qualitative results. Qualitative results are shown in Figure 4, where we compare our method with Champ [67] on the novel view animation task for the MVHumanNet dataset. Our method synthesizes temporally consistent animations, even from novel views. It highlights the advantage of leveraging human NeRF’s dense and 3D consistent appearance cue for guidance.